# SonarSplat: Novel View Synthesis of Imaging Sonar via Gaussian Splatting

### v2

@@ III-C Azimuth Streak Modeling (ASM)
Azimuth streaking, as shown in Fig. 4, is a phenomena that is frequently observed in sonar images. It is a form of saturation that occurs when the sonar receiver receives strong returns from incident angles close to parallel at a specific range bin $r_{i}$ [32]. In our model, we assign each Gaussian a probability that it will cause an azimuth streak. The benefit of defining this probability per-Gaussian is that it can be optimized to produce azimuth streaks that are multi-view consistent. Let $p_{k}$ be the probability that Gaussian $k$ will contribute to an azimuth streaking phenomena. We can then splat these per-Gaussian probabilities into a range/azimuth bin to obtain an azimuth streak probability image $P_{a}$:
$P_{a}(r_{i},\theta_{j})=\sum_{G^{\prime}_{k}\in\mathcal{G}_{i,j}}p_{k}~{}o_{k}~{}T_{k}~{}G^{\prime}_{k}(r_{i},\theta_{j})$ (9)
@@ III-C1 Adaptive Gain Module
We introduce a novel adaptive gain term, $A$, to transform the azimuth streak probability image into receiver gains. We begin designing this gain term with a few observations. First, when no Gaussian in a range bin $r_{i}$ has a high probability of azimuth streaking, the gain should be unity. Second, suppose a single Gaussian $G_{k}$ exhibits a high probability of azimuth streaking across the range bin $r_{i}$. In that case, we wish to adaptively suppress the other Gaussians in $r_{i}$ and assign a high gain to $G_{k}$. Finally, if multiple Gaussians in $r_{i}$ have high probabilities of contributing to an azimuth streak, higher gain should be assigned to those Gaussians. Following these observations, the adaptive gain term is generated for each pixel: $(r_{i},\theta_{j})$:
@@
Figure 4: We present our robotic platform and selected images from our diverse set of 8 sequences used in our experiments. We show RGB images in top right corners for visualization. Our evaluation datasets focus on smaller objects and larger-scale structures such as the test-tank basin. Azimuth streaks clearly occur in the infra_360 and piling_1 sequences. Images are range/azimuth.
@@ III-D Model Optimization
We optimize for desired parameters using gradient descent and by taking losses ($\mathcal{L}_{l1},\mathcal{L}_{ssim}$) between $\hat{I}$ and the ground truth image $I_{\text{GT}}$ [14]. Next, we note that large Gaussians cause increased training times and more distortion to the model presented in Sec. III-B. To prevent Gaussians from becoming too large, we use $\mathcal{L}_{size}=\text{ReLU}(s-s_{init})$, which encourages the scale of the Gaussians $s$ to remain smaller than their initialized scale $s_{init}$ [16]. Additionally, we find it useful for reconstruction tasks to supervise the opacities in regions with sonar returns via an opacity loss $\mathcal{L}_{o}$, similar to [35]. We compute $\mathcal{L}_{o}$ between the predicted opacity map $\hat{I}_{o}$ and $I_{\text{GT}}$ thresholded at $\tau_{o}$ to isolate strong returns. The threshold parameter is found experimentally for each dataset. $\mathcal{L}_{o}$ then becomes an $l1$ loss between $\hat{I}_{o}$ and $I^{o}_{\text{GT}}$. The final loss becomes:
@@
where $\lambda_{l1},\lambda_{o}~{}\text{and}~{}\lambda_{size}$ are the weights for the $l1$, opacity, and size losses respectively.
@@ III-E Elevation Sampling Densification Strategy (ESDS)
Splitting and cloning Gaussians in image space is unintuitive for imaging sonar given the elevation angle ambiguity. It is possible that a Gaussian can be split then immediately occluded by high opacity Gaussians in front of it. To address this, we introduce an alternative strategy for densification of Gaussians called the Elevation Sampling Densification Strategy (ESDS) shown in Fig. 3. Given predicted image $\hat{I}$ and ground truth $I_{\text{GT}}$, we sample $N_{p}$ pixels, where each pixel’s probability is proportional to the magnitude of the loss at that pixel. We initialize $N_{g}$ Gaussians on each elevation arc specified by the sampled pixel locations. This primarily serves to provide SonarSplat enough opportunities to resolve the elevation angle ambiguity and any possible occlusions by placing Gaussians at different elevation angles. As the optimization continues, low-opacity Gaussians are pruned similar to [14].
@@ III-F Implementation Details
SonarSplat is built on the gsplat repo [52]. We train our method on an NVIDIA RTX 3090 GPU with 24 GB of vRAM. We calculate the frames per second (FPS) of all methods on the same NVIDIA A6000 GPU with 50 GB of vRAM. To properly optimize for the azimuth streaking probabilities $p_{k}$ of each Gaussian, we first identify where azimuth streaks occur by calculating the average intensity of the range interval $r_{i}$. Then, for the first $N_{s}$ iterations, we train only on pixels in range bins that exceed a certain average intensity threshold $\tau_{s}$. This avoids incorrectly supervising the splat in azimuth streaked regions. Importantly, $p_{k}$ is not optimized during this interval. After $N_{s}$ iterations, we optimize only $p_{k}$ in order to capture the azimuth streaking effects. We do not optimize $\mu_{k},~{}\Sigma_{k},~{}o_{k}$, or $\nu_{k}$ during this interval. This way, we allow the Gaussians to initialize and fit the images before isolating and optimizing the azimuth streaking parameters.
@@ IV-A Robotic Platform and Datasets
We note that real, public datasets of posed sonar images suitable for novel view synthesis and reconstruction experiments do not exist. Rather, we collect sequences using a BlueRobotics BlueROV2 platform, shown in Fig. 4. Our robotic platform is equipped with a BluePrint Subsea Oculus M750-d imaging sonar operating at 1.2MHz, capturing sequences between 3-20m in range, with a 130 horizontal field of view and a 20 vertical field of view. We tilt the sonar both at 30 and 0 downwards for various sequences. We use this platform to collect 8 trajectories inspecting a range of objects and in both a test tank and river environment. We use 7 sequences for evaluating image synthesis, and 1 sequence with higher coverage of the inspected object for evaluating 3D reconstruction. Sensor poses are estimated by filtering a Doppler Velocity Logger and IMU, similar to [1, 33]. We construct ground truth for 3D reconstruction using RGB images and MASt3R, scaling the reconstruction to metric scale [17]. We find that MASt3R produces less artifacts than traditional photogrammetry pipelines like AgiSoft Metashape [21].
@@ IV-B Baselines
For baselines, we compare to Neusis [33], Neusis-NGP [50], DSC [6], and ZSplat* [35]. Since we are focused on sonar-only rendering, we compare to a modified version, ZSplat*, trained with only the sonar loss since there is no other work that has explored Gaussian splatting for imaging sonar [35]. All baselines were trained using official released code and hyperparameters were tuned to the best of our ability.
TABLE I: Novel view synthesis performance on real world held-out validation data. Note that ZSplat* is ZSplat with only the sonar loss activated for training. First, second, and third best are shown in red, orange, and yellow, respectively.
@@ IV-C Novel View Synthesis
We report validation PSNR, SSIM, and LPIPS following [14], using the lpipsPyTorch library. For our validation set, we take every 8th image from the dataset as introduced in [26]. We report SonarSplat’s performance compared to baselines in Tab. I. We find that SonarSplat consistently synthesizes more realistic sonar images, quantified by its higher PSNR and SSIM, and lower LPIPS values. Qualitative results from selected sequences are presented in Fig. 5. We also report the rendering speed in FPS of each method in Tab. I. SonarSplat and ZSplat* both leverage Gaussian splatting, which offers significantly faster rendering speeds compared to neural rendering methods. ZSplat* is optimized for high rendering speed, which will be a focus of future work.
@@ IV-D Ablation Studies
Figure 6: Qualitative comparison of SonarSplat without and with azimuth streak modeling (ASM) for piling_1. Note that if ASM is not modeled, rendered views will produce erroneous streaks (indicated with red arrows) instead of capturing the true streaks (indicated with blue arrows). Images are polar.
@@
TABLE III: Quantitative results for azimuth streak removal. We report the inverse coefficient of variation (ICV). Higher is better ($\uparrow$) and best is in bold.
@@ IV-E Azimuth Streak Removal/De-Streaking
Figure 8: Qualitative results from 3D reconstruction of SonarSplat compared to baselines. We inspect a concrete piling in a test tank. Ground truth is created using collected RGB images.
@@ IV-F Real-World 3D Reconstruction
3D reconstruction is crucial for inspection and visualization of underwater structures. In a test tank, we demonstrate inspection of a concrete piling, a structure commonly found in bridges and dams.
We show that SonarSplat learns the geometric properties of the scene by converting the optimized splat into a mesh for visualization. We compare to state-of-the-art sonar reconstruction methods: Neusis, DSC, and ZSplat* [6, 33, 35]. Qualitative results are shown in Fig. 8 and quantitative results are shown in Tab. IV. Following [33, 6], we report Root Mean Squared (RMS) $l1$ Chamfer Distance (CD-$l1$) and Hausdorff Distance (HD). We produce meshes from all methods then randomly sample 30,000 points from the surface for 30 iterations to calculate RMS. Note that neither DSC, Neusis, nor ZSplat* is able to reconstruct the object whereas SonarSplat is able to produce an accurate reconstruction of the piling.
TABLE IV: Quantitative evaluation of 3D reconstruction performance on the piling using RMS $l1$ Chamfer Distance (CD-$l1$) and Hausdorff Distance (HD). Best in bold.
Reconstruction Metrics
Method CD-$l_{1}$downarrow HD $\downarrow$
Neusis [33] 0.32 0.78
DSC [6] 0.29 0.82
ZSplat* [35] 2.93 4.88
SonarSplat (ours) 0.14 0.57
@@ V Conclusion & Future Work
We propose SonarSplat, the first sonar-only Gaussian splatting framework for imaging sonar in underwater applications. First, we adapt the sonar rendering equation for efficient range/azimuth splatting by evaluating 3D Gaussians. We model azimuth streaking within the Gaussian splatting framework, allowing us to learn per-Gaussian azimuth streaking probabilities and produce high-quality de-streaked images. We also introduce a novel densification strategy (ESDS) to enable SonarSplat to resolve elevation angle ambiguities and associated occlusions. From our extensive experiments on a diverse set of real-world data in test tank and river environments, we find that SonarSplat outperforms state-of-the-art methods in novel view synthesis (+3.2 dB PSNR) and 3D reconstruction (52% lower Chamfer Distance), and additionally allows for de-streaking of sonar images.
Future work will focus on leveraging SonarSplat for sonar data synthesis by randomizing scene parameters including acoustic reflectance and azimuth streaking probability. This can yield more diverse datasets for training sonar-based scene understanding algorithms, including multi-view target recognition algorithms [40, 38]. We will also work to improve SonarSplat’s ability to perform dense 3D reconstruction by better modeling additional acoustic phenomena like speckle noise, Lambertian scattering, and multi-path reflections. Our choice of Gaussian splatting as a 3D scene representation also opens possibilities for real-time sensor fusion and SLAM [13, 23].

### v3

@@ III-C Azimuth Streak Modeling (ASM)
Azimuth streaking, as shown in Fig. 4, is a phenomena that is frequently observed in sonar images. It is a form of saturation that occurs when the sonar receiver receives strong returns from incident angles close to 0 at a specific range bin $r_{i}$ [33]. In our model, we assign each Gaussian a probability that it will cause an azimuth streak. The benefit of defining this probability per-Gaussian is that it can be optimized to produce azimuth streaks that are multi-view consistent. Let $p_{k}$ be the probability that Gaussian $k$ will contribute to an azimuth streaking phenomena. We can then splat these per-Gaussian probabilities into a range/azimuth bin to obtain an azimuth streak probability image $P_{a}$:
$P_{a}(r_{i},\theta_{j})=\sum_{G_{k}\in\mathcal{G}_{i,j}}p_{k}~o_{k}~\overline{T}_{k}~f_{k}([r_{i},\theta_{j}])$ (11)
@@ III-C1 Adaptive Gain Module
We introduce a novel adaptive gain term, $A$, that is produced by transforming the azimuth streak probability image. We begin designing this gain term with a few observations. First, when no Gaussian in a range bin $r_{i}$ has a high probability of azimuth streaking, the gain should be unity. Second, suppose a single Gaussian $G_{k}$ exhibits a high probability of azimuth streaking across the range bin $r_{i}$. In that case, we wish to adaptively suppress the other Gaussians in $r_{i}$ and assign a high gain to $G_{k}$. Finally, if multiple Gaussians in $r_{i}$ have high probabilities of contributing to an azimuth streak, higher gain should be assigned to those Gaussians. Following these observations, the adaptive gain term is generated for each pixel: $(r_{i},\theta_{j})$:
@@
Figure 4: We present our robotic platform and selected images from our diverse set of 9 sequences used in our experiments. We show RGB images in top right corners for visualization. Our evaluation datasets focus on smaller objects and larger-scale structures such as the test-tank basin. Azimuth streaks clearly occur in the infra_360 and piling_1 sequences.
@@ III-D Model Optimization
We optimize SonarSplat using gradient descent and by taking losses ($\mathcal{L}_{l1},\mathcal{L}_{ssim}$) between $\hat{I}$ and the ground truth image $I_{\text{GT}}$ [15]. Next, we note that large Gaussians cause increased training times and more distortion to the model presented in Sec. III-B. To prevent Gaussians from becoming too large, we use $\mathcal{L}_{size}=\text{ReLU}(s-s_{init})$, which encourages the scale of the Gaussians $s$ to remain smaller than their initialized scale $s_{init}$ [38]. Additionally, we find it useful for reconstruction tasks to supervise the opacities of Gaussians via an opacity loss $\mathcal{L}_{o}$, similar to [19]. We compute $\mathcal{L}_{o}$ between the predicted opacity map $\hat{I}_{o}$ and $I_{\text{GT}}$ thresholded at $\tau_{o}$ to isolate strong returns. The threshold parameter is found experimentally for each dataset. $\mathcal{L}_{o}$ then becomes an $l1$ loss between $\hat{I}_{o}$ and $I^{o}_{\text{GT}}$. The final loss becomes:
@@
where $\lambda_{l1},\lambda_{o},~\text{and}~\lambda_{size}$ are the weights for the $l1$, opacity, and size losses respectively.
@@ III-E Elevation Sampling Densification Strategy (ESDS)
Splitting and cloning Gaussians in image space is unintuitive for imaging sonar given the elevation angle ambiguity. It is possible that a Gaussian can be split then immediately occluded by high opacity Gaussians in front of it. To address this, we introduce an alternative strategy for densification of Gaussians called the Elevation Sampling Densification Strategy (ESDS) shown in Fig. 3. Given predicted image $\hat{I}$ and ground truth $I_{\text{GT}}$, we sample $N_{p}$ pixels, where each pixel’s probability is proportional to the magnitude of the loss at that pixel. We initialize $N_{g}$ Gaussians on each elevation arc specified by the sampled pixel locations. This primarily serves to provide SonarSplat enough opportunities to resolve the elevation angle ambiguity and any possible occlusions by placing Gaussians at different elevation angles. Low-opacity Gaussians are pruned similar to [15].
@@ III-F Implementation Details
SonarSplat is built on the gsplat repo [39]. We train our method on an NVIDIA RTX 3090 GPU with 24 GB of vRAM. We calculate the frames per second (FPS) of all methods on the same NVIDIA A6000 GPU with 50 GB of vRAM. For every pixel in a sonar image with intensity above a threshold $\tau_{init}$, we initialize $N_{init}$ Gaussians on the elevation arc specified by the pixel’s range, azimuth, and the sensor’s vertical field of view. This process is repeated for all sonar images in a dataset to produce the initial set of 3D Gaussians. To properly optimize for the azimuth streaking probabilities $p_{k}$ of each Gaussian, we first identify where azimuth streaks occur by calculating the average intensity of the range interval $r_{i}$. Then, for the first $N_{s}$ iterations, we train only on pixels in range bins that exceed a certain average intensity threshold $\tau_{s}$. This avoids incorrectly supervising the splat in azimuth streaked regions. Importantly, $p_{k}$ is not optimized during this interval. After $N_{s}$ iterations, we optimize only $p_{k}$ in order to capture the azimuth streaking effects. We do not optimize $\mu_{k},~\Sigma_{k},~o_{k}$, or $\nu_{k}$ during this interval. This allows the Gaussians to fit the images before optimizing the azimuth streaking parameters.
@@ IV-A Robotic Platform and Datasets
We note that real, public datasets of posed sonar images suitable for novel view synthesis and reconstruction experiments do not exist. Rather, we collect sequences using a BlueRobotics BlueROV2 platform, shown in Fig. 4. Our robotic platform is equipped with a BluePrint Subsea Oculus M750-d imaging sonar operating at 1.2MHz, capturing sequences between 3-20m in range, with a 130 horizontal field of view and a 20 vertical field of view. The Oculus M750-d has a range resolution varying from 0.5-4 cm/pixel and angular resolution of 0.6 with a beam separation of 0.25 [40]. We tilt the sonar both at 30 and 0 downwards for various sequences. We use this platform to collect 9 trajectories inspecting a range of objects and in both a test tank and river environment. We use 7 sequences for evaluating image synthesis, and 2 sequences with higher coverage of the inspected object for evaluating 3D reconstruction. Our 3D reconstruction datasets are called Concrete Piling and Monohansett Shipwreck. Sensor poses are estimated by filtering a Doppler Velocity Logger and IMU, similar to [41, 10]. For Concrete Piling, we construct ground truth for 3D reconstruction using RGB images and MASt3R, scaling the reconstruction to metric scale [42]. For Monohansett Shipwreck, we use an underwater LiDAR scan of the shipwreck as ground truth.
@@ IV-B Baselines
For baselines, we compare to Neusis [10], Neusis-NGP [12], DSC [13], and ZSplat* [19]. Since we are focused on sonar-only rendering, we compare to a modified version, ZSplat*, trained with only the sonar loss since there is no other work that has explored Gaussian splatting for imaging sonar [19]. All baselines were trained using official code and hyperparameters were tuned to the best of our ability.
TABLE I: Novel view synthesis performance on real world held-out validation data (7 datasets). Note that ZSplat* is ZSplat with only the sonar loss activated for training. First, second, and third best are shown in red, orange, and yellow.
@@ IV-C Novel View Synthesis
We report validation PSNR, SSIM, and LPIPS following [15], using the lpipsPyTorch library. For our validation set, we take every 8th image from the dataset as introduced in [4]. We report SonarSplat’s performance compared to baselines in Tab. I. We find that SonarSplat consistently synthesizes more realistic sonar images, quantified by its higher PSNR and SSIM, and lower LPIPS values. Qualitative results from selected sequences are presented in Fig. 5. We also report the rendering speed in FPS of each method in Tab. I. SonarSplat and ZSplat* both leverage Gaussian splatting, which offers significantly faster rendering speeds compared to neural rendering methods. ZSplat* is optimized for fast rendering, which will be a focus of future work.
@@ IV-D Ablation Studies
Figure 6: Qualitative comparison of SonarSplat without and with azimuth streak modeling (ASM) for piling_1. Note that if ASM is not modeled, rendered views will produce erroneous streaks (indicated with red arrows) instead of capturing the true streaks (indicated with blue arrows).
@@
TABLE III: We report the inverse coefficient of variation (ICV) for azimuth streak removal. Best is in bold.
@@ IV-E Azimuth Streak Removal/De-Streaking
Figure 8: Qualitative results from 3D reconstruction of SonarSplat compared to baselines. We inspect Concrete Piling in a test tank. Monohansett Shipwreck was surveyed in Lake Huron, Michigan. Ground truth is created using collected RGB images and LIDAR respectively. Readers are encouraged to zoom in.
@@ IV-F Real-World 3D Reconstruction
3D reconstruction is crucial for inspection and visualization of underwater structures. In a test tank, we demonstrate inspection of Concrete Piling, a structure commonly found in bridges and dams. We also surveyed Monohansett Shipwreck, a wooden shipwreck at a depth of 18 feet in Lake Huron, Michigan, measuring 160 feet in length and 30 feet in width.
We show that SonarSplat learns the geometric properties of the scene by converting the optimized splat into a mesh for visualization. We sample a set of points from the probability density functions described by the Gaussians, then query the opacities and use marching cubes to produce a mesh. Specifically, we used between 10-50 samples per Gaussian depending on the size of the scene and the density of the Gaussians. We compare to state-of-the-art sonar reconstruction methods: Neusis, DSC, and ZSplat* [13, 10, 19]. Additionally, we ablate ESDS (denoted as “SonarSplat w/o ESDS”). Qualitative results are shown in Fig. 8 and quantitative results are shown in Tab. IV. Following [10, 13], we report Root Mean Squared (RMS) $l1$ Chamfer Distance (CD-$l1$) and Hausdorff Distance (HD), which are both in meters. We produce meshes from all methods then randomly sample 30,000 points from the surface for 30 iterations to calculate RMS. We use ICP to align the predictions and ground truth meshes then crop the predictions to the bounding box of the ground truth. Note that neither DSC, Neusis, nor ZSplat* is able to reconstruct the object whereas SonarSplat is able to produce an accurate reconstruction of the piling. For Monohansett Shipwreck, we see that SonarSplat captures details like the wooden ribs of the shipwreck more accurately than baselines.
TABLE IV: Quantitative evaluation of 3D reconstruction using RMS $l1$ Chamfer Distance (CD-$l1$) and Hausdorff Distance (HD). Units are meters. Best in bold.
3D Reconstruction Performance
Concrete Piling Monohansett Shipwreck
Method CD-$l_{1}$ $\downarrow$ HD $\downarrow$ CD-$l_{1}$ $\downarrow$ HD $\downarrow$
Neusis [10] 0.32 0.78 1.75 3.67
DSC [13] 0.29 0.82 1.60 6.77
ZSplat* [19] 2.93 4.88 3.06 8.30
SonarSplat w/o ESDS 0.92 1.77 3.62 12.52
SonarSplat (ours) 0.14 0.57 0.29 1.06
@@ V Conclusion & Future Work
We propose SonarSplat, the first sonar-only Gaussian splatting framework for imaging sonar in underwater applications. First, we adapt the sonar rendering equation for efficient range/azimuth splatting by evaluating 3D Gaussians. We model azimuth streaking within the Gaussian splatting framework, allowing us to learn per-Gaussian azimuth streaking probabilities and produce high-quality de-streaked images. We also introduce a novel densification strategy (ESDS) to enable SonarSplat to resolve elevation angle ambiguities and associated occlusions. From our extensive experiments on a diverse set of real-world data in test tank and lake environments, we find that SonarSplat outperforms state-of-the-art methods in novel view synthesis (+3.2 dB PSNR) and 3D reconstruction (77% lower Chamfer Distance), and additionally allows for de-streaking of sonar images.
Future work will focus on leveraging SonarSplat for sonar data synthesis by randomizing scene parameters including acoustic reflectance and azimuth streaking probability. This can yield more diverse datasets for training sonar-based scene understanding algorithms, including multi-view target recognition algorithms [46, 47]. We will also work to improve SonarSplat’s ability to perform dense 3D reconstruction by better modeling additional acoustic phenomena like speckle noise and multi-path reflections. Our choice of Gaussian splatting as a 3D scene representation also opens possibilities for real-time sensor fusion and SLAM [24].